OpenStudy (anonymous):

(5/4x^2y) - (y/14xz)

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

\[\frac{ 5 }{ 4x ^{2}y }-\frac{ y }{ 14xz }=\frac{ 35z }{ 28x ^{2}yz }-\frac{ 2xy ^{2} }{ 28x^{2}yz }\]\[=\frac{ 35z-2xy ^{2} }{ 28^{2}yz }\]